### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized code generation but are require significant resources and tend to over-generalize, limiting their task-specific efficiency. Fine-tuning smaller, open-source LLMs is a cost-effective alternative, yet standard supervised approaches rely solely on correct examples, overlooking valuable insights from failures. We introduce CodeLutra, a new framework that leverages both correct and incorrect code attempts. Instead of purely instructing with correct solutions, CodeLutra uses iterative preference-based refinement, comparing successful and failed outputs to better approximate desired results. This process narrows the performance gap with state-of-the-art, larger models, without requiring massive datasets or auxiliary models. For example, on a challenging data science coding task, using only 500 samples improved Llama-3-8B’s accuracy from 28.2% to 48.6%, approaching GPT-4’s level. By capitalizing on both successes and mistakes, CodeLutra offers a scalable, efficient path to high-quality code generation, making smaller open-source models more competitive with leading closed-source alternatives.
[Uncaptioned image]CodeLutra: Boosting LLM Code Generation via Preference-Guided Refinement
@@ 1 Introduction
Large language models (LLMs) have revolutionized numerous domains, consistently delivering great performance across different tasks (Brown et al., 2020; Achiam et al., 2023; Anthropic, 2023; OpenAI, 2023). Among these applications, code generation stands out as particularly promising. Models pre-trained on extensive code repositories have demonstrated an impressive capability to solve diverse programming challenges (Li et al., 2022; Zheng et al., 2023; Chen et al., 2021b; Wang et al., 2023b).
@@
To address this challenge, we introduce CodeLutra, a framework that iteratively improves the performance of a given LLM without relying on vast external datasets or larger auxiliary models. Unlike traditional fine-tuning methods that rely exclusively on correct code solutions, CodeLutra leverages both successful and failed code generated by the model itself, creating self-generated comparative data. While prior works explored using execution results to enhance code generation, they often require extensive datasets to train reward models for feedback Le et al. (2022); Shojaee et al. (2023) or rely on prompt engineering with ultra-large LLMs Chen et al. (2023); Ni et al. (2024). In contrast, our work collects successful and failed attempts and enables models to iteratively learn, distinguish, and improve from these examples. Our approach demonstrates that even with limited data at hand, substantial gains in code generation quality can be achieved, closing the gap between smaller fine-tuned models and the top-tier LLMs. To harness the comparative data, a key component of CodeLutra is the preference-guided refinement mechanism that compares correct and incorrect code snippets, iteratively refining the model’s understanding of code quality. With each iteration, the model generates code solutions, evaluates their correctness, and updates its parameters based on the evolving preference dataset. This process allows for continuous improvements of the base model, making the framework effective even with limited initial data (e.g., a few hundred samples). The failed code attempts are invaluable for refining models, since they provide concrete examples of common errors and enable the model to learn strategies for avoiding similar mistakes in future generations.
We comprehensively evaluate the effectiveness of CodeLutra on challenging data query and data science tasks, where the LLM is tasked with generating the correct SQL or Python code to solve a given problem. We compare CodeLutra with 13 open-source and closed-source LLMs that are competitive in code generation. Notably, on the data query task, our framework allows Llama-3-8B (Dubey et al., 2024) to achieve an execution accuracy of 76.6%, which exceeds GPT-4’s 74.4%. Under a challenging data science task, we find that using just 500 samples improved Llama-3-8B from an accuracy of 28.2% to 48.6%, approaching the performance of GPT-4. This demonstrates that CodeLutra achieves strong results even with a limited number of high-quality annotations. Moreover, we observe the consistent performance gains of CodeLutra on different base models, including Gemma-7B (Team et al., 2024) and StarCoder-7B (Dai and Kumar, 2023). These findings highlight the potential of CodeLutra in closing the performance gap between open-source and closed-source models. To summarize our key contributions:
@@
    We present CodeLutra, a novel framework that iteratively improves LLMs for code generation using self-generated comparative data from both successful and failed code attempts, enabling low-performing models to rival top-tier solutions without external datasets or ultra-large LLMs’ feedback.
@@ 2 Related Work
Preference learning guides language models toward outputs aligned with human preferences. Approaches like RLHF (Ziegler et al., 2019; Ouyang et al., 2022) effectively align LLMs but face inefficiencies and hyperparameter sensitivity (Christiano et al., 2017; Bai et al., 2022). Recent works focus on closed-form loss functions using offline preference data, such as DPO (Rafailov et al., 2023) and its extensions (Liu et al., 2023; Ethayarajh et al., 2024). Iterative DPO, where models generate their own preference data, shows strong performance in the alignment and math reasoning tasks (Liu et al., 2024; Xiong et al., 2024; Yuan et al., 2024; Pang et al., 2024). Different from prior works, we present the first comprehensive study on iterative preference optimization for code generation, achieving GPT-4-level results and paving the way for future advancements.
@@
LLMs trained on large code corpora excel in tasks like code generation (Chen et al., 2021c; Zhang et al., 2022), program repair (Xia and Zhang, 2022), and software testing (Chen et al., 2023). Models like WizardCoder (Luo et al., 2023) and DS-Coder (Li et al., 2023) enhance these capabilities through repository-level organization and retrieval-augmented techniques (Borzunov et al., 2024). Instruction fine-tuning, as in CodeInstruct (Wang et al., 2024b), further improves alignment with human coding preferences. Different from other methods that also rely on program execution results to help model training (Le et al., 2022; Shojaee et al., 2023; Ni et al., 2024; Zhang et al., 2024), which relies on reward model or prompt engineering, we iteratively refine small LLMs using self-generated preference data, avoiding external datasets and larger models. Additionally, we unify SFT and preference learning with a dual-loss approach, maximizing the likelihood of correct solutions in one stage. We include an expansive literature discussion in Appendix A.
@@ 3 Preliminaries
Code generation demands not just syntactical correctness but also a deep understanding of logical and domain-specific nuances, which complicates model training. A major limitation of SFT is that it solely maximizes the likelihood of providing correct code, which restricts the model’s ability to learn from its own mistakes. Since the training process focuses exclusively on provided examples, the LLM doesn’t receive the gradient from incorrect or suboptimal code. For instance, if the model only predicts wrongly in the final token in a code snippet, the overall probability $P({y}|{x})$ might still remain high as the preceding tokens are correct. Correspondingly, the SFT loss is very small in that case. However, this single erroneous token can render the entire code nonfunctional or introduce subtle bugs, significantly affecting execution correctness despite a high likelihood score, which is different from general NLP tasks. This reliance on only correct examples limits the model’s capacity to identify and recover from errors, reducing its ability to handle more complex or nuanced coding tasks. This motivates our framework CodeLutra, which leverages both successful and failed code generation attempts.
@@ 4 CodeLutra
In this section, we introduce CodeLutra, a framework designed to comparatively learn from both correct and incorrect code generations. CodeLutra delivers substantial performance gains, achieving performance comparable to more advanced models like GPT-4 (OpenAI, 2023), even with limited initial data. The pseudo-code is provided in the Appendix B.1.
@@
While one can directly employ a preference optimization approach like DPO (Rafailov et al., 2023) on our curated dataset $\mathcal{D}_{t}$, this approach presents a notable limitation due to its tendency to decrease the likelihood of both correct and rejected code during training. This is evidenced by the dashed lines in Figure 2, which is also observed in Pal et al. (2024); Feng et al. (2024). This diminishing likelihood can significantly impact our framework, especially since our correct code are critical for successfully solving the assigned tasks. It is crucial, therefore, to prioritize the likelihood of correct solutions to the task at hand.
@@
To address this limitation, we employ a dual-loss function integrating DPO with an additional term, which regularizes the training to prevent decreasing likelihood Liu et al. (2024). Specifically, we maximize the likelihood of correct solutions in dataset $\mathcal{D}_{t}^{c}=\{\left(x_{i},\hat{y}_{i}^{c_{k}}\right)\mid\text{ for all}x_{i}\in\mathcal{D}\text{ and }k\in[K]\}$. The overall loss function is defined as:
$\pi_{t+1}=\operatorname{argmin}_{\pi_{\theta}}\Bigg{[}$ $-\mathbb{E}\Bigg{[}\log\sigma\Bigg{(}\beta\Bigg{(}\log\frac{\pi_{\theta}(y_{i}^{c}|x_{i})}{\pi_{\text{t}}(y_{i}^{c}|x_{i})}$
$\quad-\log\frac{\pi_{\theta}(y_{i}^{r}|x_{i})}{\pi_{\text{t}}(y_{i}^{r}|x_{i})}\Bigg{)}\Bigg{)}\Bigg{]}$
$-\lambda\mathbb{E}\Big{[}\log\pi_{\theta}(y_{i}^{c}|x_{i})\Big{]}\Bigg{]},$ (2)
@@ 5.1 Experimental Setup
To evaluate the effectiveness of our proposed framework, we perform experiments on two tasks: Data Query and Data Science. Both tasks reflect common and practical challenges in fields such as business, healthcare, and scientific computing, where precise code generation is critical for solving data-related problems while receiving limited attention. We provide more details and examples of the two tasks in the Appendix B.4.
@@
We conduct our experiments on two cross-domain datasets for data query, Spider (Yu et al., 2018) and BIRD (Wang et al., 2023a), and a data science dataset, DS-1000 (Lai et al., 2023). We provide more details in the Appendix B.3.
@@
To evaluate the effectiveness of our method, we compare it against three categories of baselines. For a fair comparison, all reported results are based on the same prompt. We report the details of baseline setup in Appendix B.6.
  • Open-source LLMs: We benchmark our method against competitive open-source LLMs, including models pre-trained on general datasets such as Llama-3-8B (Dubey et al., 2024), Gemma-7B (Team et al., 2024), and Llama-3-70B-Instruct (Dubey et al., 2024). Additionally, we compare against LLMs pre-trained specifically on coding datasets, such as Codellama-7B (Xu and Zhang, 2023), StarCoder-7B (Lozhkov et al., 2024), and Codestral-22B (Brown and Lee, 2023).
@@
For main results, we apply our framework to the Llama-3-8B base model (Dubey et al., 2024), denoted as $\pi_{0}$ (see Section 5.2 for more backbone results). We use a zero-shot prompt containing the question along with reference information. For different answer collections, we employ the best-of-$n$ strategy by sampling 16 responses at the temperature of 1.0. We train 1 epoch per iteration and perform four iterations in total, resulting in models $\{\pi_{1},\pi_{2},\pi_{3},\pi_{4}\}$. These models are evaluated as described in the following sections. For more experimental details, please refer to the Appendix B.2.

### v3

@@ Abstract
Large Language Models (LLMs) have revolutionized code generation but require significant resources and tend to over-generalize, limiting their task-specific efficiency. Fine-tuning smaller, open-source LLMs is a cost-effective alternative, yet standard supervised approaches rely solely on correct examples, overlooking valuable insights from failures. We explore recent developments in preference-based post-training to code generation, leveraging both correct and incorrect code attempts. Instead of solely relying on correct examples, our framework CodeLutra iteratively refines the model by comparing successful and unsuccessful outputs, thereby more accurately aligning the generated code with desired outcomes. This process narrows the performance gap with state-of-the-art, larger models, without requiring massive datasets or auxiliary models. For example, on a challenging data science coding task, using only 500 samples improved Llama-3-8B’s accuracy from 28.2% to 48.6%, approaching GPT-4’s level. CodeLutra provides a scalable and efficient strategy for high-quality code generation, helping smaller open-source models narrow the performance gap with state-of-the-art closed-source systems.
@@ 1 Introduction
Large language models (LLMs) have revolutionized numerous domains, consistently delivering great performance across different tasks (Brown et al., 2020; Achiam et al., 2023; Anthropic, 2023; OpenAI, 2023). Among these applications, code generation stands out as particularly promising. Models pre-trained on extensive code repositories have demonstrated an impressive capability to solve diverse programming challenges (Li et al., 2022; Zheng et al., 2023; Chen et al., 2021b; Wang et al., 2023).
@@
To address this challenge, we introduce CodeLutra, a framework that iteratively improves the performance of a given LLM without relying on vast external datasets or larger auxiliary models. Unlike traditional fine-tuning methods that rely exclusively on correct code solutions, our framework CodeLutra leverages both successful and failed code generated by the model itself, creating self-generated comparative data. While prior works explored using execution results to enhance code generation, they often require extensive datasets to train reward models for feedback (Le et al., 2022; Shojaee et al., 2023) or rely on prompt engineering with ultra-large LLMs (Chen et al., 2023; Ni et al., 2024). In contrast, our work collects successful and failed attempts and enables models to iteratively learn, distinguish, and improve from these examples. Our approach demonstrates that even with limited data at hand, substantial gains in code generation quality can be achieved, closing the gap between smaller fine-tuned models and the top-tier LLMs. To harness the comparative data, a key component of CodeLutra is the preference-guided refinement mechanism that compares correct and incorrect code snippets, iteratively refining the model’s understanding of code quality. With each iteration, the model generates code solutions, evaluates their correctness, and updates its parameters based on the evolving preference dataset. This process allows for continuous improvements of the base model, making the framework effective even with limited initial data (e.g., a few hundred samples). The failed code attempts are invaluable for refining models, since they provide concrete examples of common errors and enable the model to learn strategies for avoiding similar mistakes in future generations.
We comprehensively evaluate the effectiveness of CodeLutra on challenging data query and data science tasks, where the LLM is tasked with generating the correct SQL or Python code to solve a given problem. We compare CodeLutra with 13 open-source and closed-source LLMs that are competitive in code generation. Notably, on the data query task, our framework allows Llama-3-8B (Dubey et al., 2024) to achieve an execution accuracy of 76.6%, which exceeds GPT-4’s 74.4%. Under a challenging data science task, we find that using just 500 samples improved Llama-3-8B from an accuracy of 28.2% to 48.6%, approaching the performance of GPT-4. This demonstrates that CodeLutra achieves strong results even with a limited number of high-quality annotations. Moreover, we observe the consistent performance gains of CodeLutra on different base models, including Gemma-7B (Team et al., 2024) and StarCoder-7B (Dai & Kumar, 2023). These findings highlight the potential of CodeLutra in closing the performance gap between open-source and closed-source models. To summarize our key contributions:
@@
    We explore preference-based learning for code generation through CodeLutra, which leverages self-generated comparative data from both successful and failed code attempts, enabling low-performing models to rival top-tier solutions without external datasets or ultra-large LLMs’ feedback.
@@ 2 Related Work
Preference learning guides language models toward outputs aligned with human preferences. Approaches like RLHF (Ziegler et al., 2019; Ouyang et al., 2022) effectively align LLMs but face inefficiencies and hyperparameter sensitivity (Christiano et al., 2017; Bai et al., 2022). Recent works focus on closed-form loss functions using offline preference data, such as DPO (Rafailov et al., 2023) and its extensions (Liu et al., 2023b; Ethayarajh et al., 2024). Iterative DPO, where models generate their own preference data, shows strong performance in the alignment and math reasoning tasks (Liu et al., 2024; Xiong et al., 2024; Yuan et al., 2024; Pang et al., 2024). Our work represents the first comprehensive study on iterative preference optimization specifically tailored for code generation, achieving GPT-4-level results with significantly fewer resources. While previous approaches (Le et al., 2022; Shojaee et al., 2023; Ni et al., 2024; Chen et al., 2024) leverage execution results predominantly to refine incorrect code until a correct solution emerges, they often rely heavily on complex reward signals from critic models (Le et al., 2022; Shojaee et al., 2023) or sophisticated prompt engineering (Chen et al., 2024; Ni et al., 2024). The key novelty of our method lies in using execution outcomes to probabilistically differentiate correct from incorrect generations via a Bradley–Terry model, representing a fundamental conceptual shift from existing approaches. This conceptual shift allows our framework to remain simpler, more generalizable, and less dependent on proprietary large-scale models: rather than crafting intricate instructions tailored to advanced architectures or relying on external ground truths like unit tests, we directly let the model learn from failures by contrasting them with successes, thereby improving its coding proficiency over iterative training cycles. Compared to other methods, which introduce complexities that may not scale down to smaller models or open-source environments, our minimalistic strategy consistently demonstrates strong performance.
@@
LLMs trained on large code corpora excel in tasks like code generation (Chen et al., 2021c; Zhang et al., 2022), program repair (Xia & Zhang, 2022), and software testing (Chen et al., 2023). Models like WizardCoder (Luo et al., 2023) and DS-Coder (Li et al., 2023b) enhance these capabilities through repository-level organization and retrieval-augmented techniques (Lu et al., 2022). Instruction fine-tuning, as in CodeInstruct (Li et al., 2023a), further improves alignment with human coding preferences. Unlike existing methods that rely on program execution results (Le et al., 2022; Shojaee et al., 2023; Ni et al., 2024; Zhang et al., 2024), which depend on complex reward models or elaborate prompt engineering, our novel approach iteratively refines small LLMs using self-generated preference data, completely eliminating the need for external datasets and larger models. Additionally, we employ a unification of SFT and preference learning through our dual-loss approach, maximizing the likelihood of correct solutions in one stage. Furthermore, although theoretical insights into dual loss objectives have been discussed in other work (Liu et al., 2024), our approach distinctively refines and applies these principles to the uniquely challenging domain of code generation—showing that the integration of execution feedback into a direct, probabilistic learning framework can yield practical gains without the pitfalls of overly elaborate interventions. While prior iterative reasoning methods excel at mathematical tasks (Pang et al., 2024), code generation poses distinct challenges due to strict syntactic and semantic correctness requirements. In mathematical reasoning, a correct final output does not necessarily guarantee an entirely accurate answer (Wang et al., 2024). However, in code generation, even minor errors in the code can result in completely incorrect execution outcomes. The novelty of our proposed framework, CodeLutra, lies in its unique integration of both successful and failed attempts using a Bradley–Terry model, enabling the refinement of outputs without any reliance on explicit feedback or ground-truth labels. By generalizing from errors and iteratively improving attempts, CodeLutra consistently reduces syntactic and logical flaws while maintaining the quality of correct solutions. Empirically, it matches or even surpasses advanced models like GPT-4, advancing the democratization of high-quality code generation in a domain with strict correctness requirements.
@@ 3 Preliminaries
LLMs for code generation.
LLMs are pre-trained on diverse datasets encompassing both natural and programming languages. In code generation tasks, an LLM receives a prompt—such as a natural language description, and generates the corresponding code by predicting the next token in the sequence. Formally, code generation is modeled as the conditional probability of a code sequence ${y}=(y_{1},y_{2},\dots,y_{T})$ given an input prompt ${x}$:
$P({y}|{x})=\prod_{t=1}^{T}P(y_{t}|y_{<t},{x}),$ (1)
where ${x}$ is the input prompt, ${y}$ is the generated code sequence of length $T$, and $y_{<t}=(y_{1},y_{2},\dots,y_{t-1})$ represents the tokens generated before time step $t$.
@@
Code generation demands not just syntactical correctness but also a deep understanding of logical and domain-specific nuances, which complicates model training. A major limitation of SFT is that it solely maximizes the likelihood of providing correct code, which restricts the model’s ability to learn from its own mistakes. Since the training process focuses exclusively on provided examples, the LLM doesn’t receive the gradient from incorrect or suboptimal code. For instance, if the model only predicts wrongly in the final token in a code snippet, the overall probability $P({y}|{x})$ might still remain high as the preceding tokens are correct. Correspondingly, the SFT loss is very small in that case. However, unlike natural language generation where minor errors might still preserve meaning, in code generation, this single erroneous token can render the entire code nonfunctional or introduce subtle bugs, significantly affecting execution correctness despite a high likelihood score. This binary nature of code correctness—where code either executes perfectly or fails completely—creates a unique challenge that standard SFT approaches struggle to address. This reliance on only correct examples limits the model’s capacity to identify and recover from errors, reducing its ability to handle more complex or nuanced coding tasks. This motivates our framework CodeLutra, which leverages both successful and failed code generation attempts.
@@ 4 CodeLutra
In this section, we introduce CodeLutra, a framework designed to comparatively learn from both correct and incorrect code generations. CodeLutra delivers substantial performance gains, achieving performance comparable to more advanced models like GPT-4 (OpenAI, 2023), even with limited initial data. We summarize our algorithm in implementation in the Algorithm 1.
@@
While one can directly employ a preference optimization approach like DPO (Rafailov et al., 2023) on our curated dataset $\mathcal{D}_{t}$, this approach presents a notable limitation due to its tendency to decrease the likelihood of both correct and rejected code during training. This is evidenced by the dashed lines in Figure 2, which is also observed in Pal et al. (2024); Feng et al. (2024); Liu et al. (2024). This occurs because DPO relies on proxy rewards from limited data, leading to misdirected gradient updates that cause the model to conservatively reduce likelihood for both response types (Liu et al., 2024). This diminishing likelihood can significantly impact our framework, especially since our correct code are critical for successfully solving the assigned tasks. It is crucial, therefore, to prioritize the likelihood of correct solutions to the task at hand.
To address this limitation, we employ a dual-loss function integrating DPO with an additional term, which regularizes the training to prevent decreasing likelihood (Liu et al., 2024). Specifically, we maximize the likelihood of correct solutions in dataset $\mathcal{D}_{t}^{c}=\{\left(x_{i},\hat{y}_{i}^{c_{k}}\right)\mid\text{ for all}x_{i}\in\mathcal{D}\text{ and }k\in[K]\}$. The overall loss function is defined as:
@@
Algorithm 1 CodeLutra
1:Training set $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$; initial base model $\pi_{0}$; responses per input $M$; preference pairs $K$; iterations $T$; hyper-parameter $\lambda$
2:Fine-tuned model $\pi_{T}$
3:for $t\leftarrow 0$ to $T-1$ do
4:     Initialize preference dataset $\mathcal{D}_{t}\leftarrow\emptyset$
5:     for all $x_{i}\in\mathcal{D}$ do
6:         Initialize chosen code set $Y_{i}^{(c)}\leftarrow\emptyset$
7:         Initialize rejected code set $Y_{i}^{(r)}\leftarrow\emptyset$
8:         for $k\leftarrow 1$ to $M$ do
9:              Generate response $\hat{y}_{i}^{k}\sim\pi_{t}(x_{i})$
10:              if execution result of $\hat{y}_{i}^{k}$ matches $y_{i}$ then
11:                  Add $\hat{y}_{i}^{k}$ to $Y_{i}^{(c)}$
12:              else
13:                  Add $\hat{y}_{i}^{k}$ to $Y_{i}^{(r)}$
14:              end if
15:         end for
16:         for $k\leftarrow 1$ to $K$ do
17:              Randomly sample $\hat{y}_{i}^{c_{k}}$ from $Y_{i}^{(c)}$ (with replacement if $|Y_{i}^{(c)}|<K$)
18:              Randomly sample $\hat{y}_{i}^{r_{k}}$ from $Y_{i}^{(r)}$ (with replacement if $|Y_{i}^{(r)}|<K$)
19:              Add $(x_{i},\hat{y}_{i}^{c_{k}},\hat{y}_{i}^{r_{k}})$ to $\mathcal{D}_{t}$
20:         end for
21:     end for
22:     Update model $\pi_{t+1}$ by minimizing the combined loss:
23:end for
@@ 5.1 Experimental Setup
To evaluate the effectiveness of our proposed framework, we perform experiments on two tasks: Data Query and Data Science. Both tasks reflect common and practical challenges in fields such as business, healthcare, and scientific computing, where precise code generation is critical for solving data-related problems while receiving limited attention. We provide more details and examples of the two tasks in the Appendix A.3.
@@
We conduct our experiments on two cross-domain datasets for data query, Spider (Yu et al., 2018) and BIRD (Li et al., 2024), and a data science dataset, DS-1000 (Lai et al., 2023). We provide more details in the Appendix A.2.
@@
To evaluate the effectiveness of our method, we compare it against three categories of baselines. For a fair comparison, all reported results are based on the same prompt. We report the details of baseline setup in Appendix A.6.
  • Open-source LLMs: We benchmark our method against competitive open-source LLMs, including models pre-trained on general datasets such as Llama-3-8B (Dubey et al., 2024), Gemma-7B (Team et al., 2024), and Llama-3-70B-Instruct (Dubey et al., 2024). Additionally, we compare against LLMs pre-trained specifically on coding datasets, such as Codellama-7B (Roziere et al., 2023), StarCoder-7B (Lozhkov et al., 2024), and Codestral-22B (Mistral AI, 2024).
@@
For main results, we apply our framework to the Llama-3-8B base model (Dubey et al., 2024), denoted as $\pi_{0}$ (see Section 5.2 for more backbone results). We use a zero-shot prompt containing the question along with reference information. For different answer collections, we employ the best-of-$n$ strategy by sampling 10 responses at the temperature of 1.0. We train 1 epoch per iteration and perform four iterations in total, resulting in models $\{\pi_{1},\pi_{2},\pi_{3},\pi_{4}\}$. For verification of code correctness, we employ task-specific approaches: For data query tasks, we execute the generated SQL queries against the same database schema as the reference solutions and directly compare execution results, ensuring functional equivalence regardless of syntactic differences. For data science tasks, we leverage the DS-1000 evaluation framework with standardized test cases that verify functional equivalence between generated and reference code. These test cases comprehensively cover edge cases and typical usage patterns to ensure thorough validation. Our correctness verification process is fully automated, allowing efficient processing of thousands of code samples during both training and evaluation phases. For more experimental details, please refer to the Appendix A.1.